An investor longs a 165 strike Mar 20, 2020 call on HON with a premium of 12.75 and shorts a 180 strike Mar 20, 2020 call on HON with a premium of 4.40. The quoted risk free rate is 1.72% on a discount basis.

What is the breakeven price?
Answer #1

Profit=MAX(S-165,0)-12.75-MAX(S-180,0)+4.40

180>S>165: Profit=S-165-12.75-0+4.40=S-173.35

180<S: Profit=S-165-12.75-S+180+4.40=6.65000

S<165: Profit=-12.75+4.40=-8.35000

Breakeven stock price=173.35